OpenStudy (anonymous):

Which equation represents y=-x^2+6x+7 in vertex form?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

Use completion of square method: y=x2−6x+32−32+7 y=(x−3)2−2

OpenStudy (anonymous):

your vertex is (3,-2)!

OpenStudy (anonymous):

your vertex is (3,-2)!

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

any more?